# ACING: Actor-Critic for Instruction Learning in Black-Box Large Language Models

### v1

@@ 3.2 Actor-Critic Continuum Bandits
^$\hat{\nabla}_{\theta}J_{Q}(\mathbf{w})=\nabla_{\mathbf{w}}Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)\left(Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)-r\left(\mathbf{a}_{t}\right)\right),$ (5)
where $Q$ is the critic-network with parameters $\mathbf{w}$. The actor-critic learning process alternates between policy-network and critic-network updates.
3.3 Enhancing Exploration with Entropy
In general, RL learning approaches aim to maximize the expected cumulative rewards up to a horizon $T$, expressed as $\max_{\pi(\cdot)}\sum_{t=1}^{T}\mathbb{E}_{a_{t}\sim\pi(\cdot)}\left[r(a_{t})\right].$ In contrast, black-box prompt learning, framed as a best-arm identification problem, seeks the best prompt without necessarily maximizing the expected cumulative rewards. To maximize the cumulative rewards, the agent must manage the exploration-exploitation trade-off. Conversely, for the best-arm identification problem, the agent should explore as extensively as possible within the $T$ evaluation budget to potentially identify the best action, leading to a pure exploration setting. Although it might seem that the agent should act purely at random to find the best action, some degree of exploitation may still be beneficial to guide the agent toward the optimal choice.
To achieve a high exploration rate while exploiting prior data, we consider the maximum entropy objective, as discussed in previous RL works Ziebart (2010); Haarnoja et al. (2018). This objective promotes stochastic policies by augmenting the reward with an entropy term, expressed as:
@@ 3.3 Enhancing Exploration with Entropy
where the temperature $\alpha$ controls the stochasticity of the agent. The above objective is a simplification of the objective in the Soft Actor-Critic (SAC) algorithm Haarnoja et al. (2018), eliminating the state, as we consider a stateless environment and focus on instantaneous rather than cumulative rewards. The term $-\mathbb{E}_{a\sim\pi(.;\theta)}\left[\log\pi(a)\right]$ represents the entropy of the policy, which we seek to maximize alongside the critic values, as entropy maximization is crucial for encouraging the actor to explore different actions. A high $\alpha$ encourages exploration, whereas a lower value of $\alpha$ favors exploitation.
Furthermore, to avoid fine-tuning, and to achieve adaptive exploration, the entropy coefficient $\alpha$ can be adjusted dynamically to maintain a target entropy level Haarnoja et al. (2019), $H_{\text{target}}$, ensuring sufficient exploration, as follows:
@@
We use a stochastic approximation of gradients, specifically the Adam algorithm Kingma and Ba (2017); Reddi et al. (2018), to minimize the losses via stochastic gradient descent. The learning rates $\lambda$, $\beta$, and $\gamma$ are used for the Q-network, the policy networks, and the temperature, respectively.
Figure 2: Illustration of the prompt generation and testing inside the environment.
4 Methodology
The methodology of our proposed algorithm, ACING, is illustrated in Fig. 1, showing the actor-critic interaction with the environment. In Fig. 2, the environment is zoomed in, using the larger_animal dataset as an example. In the following, we provide a detailed explanation of the full methodology.
Overview. In each iteration $t\leq T$, the actor-critic agent generates a continuous vector “action” $\mathbf{a}$ (step 1). This action is then played on the environment, which projects $\mathbf{a}$ into the appropriate space using a fixed projection matrix $P$ to obtain $\mathbf{z}$. The environment then concatenates the projected vector $\mathbf{z}$ with a set of exemplars’ embeddings from $\mathcal{E}$ and feeds it into a white-box LLM $h$ (step 2). The white-box LLM produces a discrete prompt, $\tau$, which is evaluated using the validation dataset $\mathcal{V}$ based on the responses from the black-box LLM $f$ (step 3). The black-box LLM generates a prediction, which is then contrasted to the true labels of the validation examples, and a score function provides a final reward to the critic. This reward is used to compute the critic’s loss and update both the critic and actor networks accordingly.
@@ 4 Methodology
By repeating steps 1 to 4 until the allowed budget $T$ is exhausted, the agent returns the best instruction prompt, denoted as $\tau^{\star}$. This prompt is then evaluated using the test set $\mathcal{T}$ on the black-box LLM, yielding the final score $\mathcal{S}^{\star}=\frac{1}{m^{\prime}}\sum_{i=1}^{m^{\prime}}q(f(\tau^{\star},x^{\prime}_{i}),y^{\prime}_{i})$ (using the evaluation template depicted in Fig. 1 bottom right).
Category Task Human Instruction ACING Instruction (Ours) Human Score Our Score
Spelling Second_word_letter Extract the second letter of the input word. Output the first letter of each word that is a vowel (a, e, i, o, u) Example: Input: year Output: e Input: trust Output: r Input: qualification Output: u Input: defendant Output: e 0.96 0.59
Syntax Negation Negate the input sentence. Flip the truth value of the statements in the input 0.81 0.82
Lexical Semantics Antonyms Write a word that means the opposite of the input word. Take a word and change it to its opposite 0.70 0.82
Synonyms Write a word with a similar meaning to the input word. Input a word that is a synonym for the word that was output 0.14 0.13
Phonetics Rhymes Write a word that rhymes with the input word. Input the word that the program thought I was inputting and then output the word that program thought I was inputting 0.61 1.00
Semantics Cause_and_effect Find which of the two given cause and effect sentences is the cause. Find the sentence that is the cause of the effect in the pair of sentences 0.97 0.92
Style Informal_to_formal Rephrase the sentence in formal language. Convert the input into output using the same word order and with the same meaning 0.63 0.5
Multi- lingual Translation_en-de Translate the word into German. Provide a translation for each word in the English text into German 0.81 0.84
Translation_en-es Translate the word into Spanish. Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 0.89 0.88
Translation_en-fr Translate the word into French. Create a program that would take an English word as input and output its French equivalent 0.86 0.87
GLUE Sentiment Determine whether a movie review is positive or negative. Classify each input as positive or negative based on the assessment of the corresponding movie 0.89 0.90
Sentence_similarity Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly. Find a sentence pair that is probably not similar, and the output is 3 - probably 0.00 0.21
median score 0.81 0.83
average score 0.69 0.71
@@
Table 1: Tasks from the instruction-induction datasets where the human and ACING test scores differed. For each task, we provide the corresponding human instruction as proposed in Honovich et al. (2023) and our best discovered instruction. We tested these instructions on the test dataset and report their respective scores.
Category Task APE InstructZero INSTINCT ACING
Spelling Letters_list 0.59 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00)
First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Second_word_letter 0.00 (0.00) 0.35 (0.09) 0.39 (0.28) 0.70 (0.15)
Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.95 (0.03)
Active_to_passive 1.00 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00)
Negation 0.79 (0.00) 0.65 (0.10) 0.58 (0.22) 0.71 (0.06)
Lexical Semantics Antonyms 0.79 (0.02) 0.76 (0.00) 0.84 (0.01) 0.74 (0.01)
Synonyms 0.14 (0.01) 0.22 (0.11) 0.19 (0.08) 0.13 (0.02)
Word_unscrambling 0.54 (0.00) 0.59 (0.06) 0.54 (0.02) 0.50 (0.07)
Phonetics Rhymes 0.59 (0.01) 0.99 (0.01) 0.36 (0.04) 0.57 (0.31)
Numerical Sum 0.87 (0.01) 1.00 (0.00) 0.70 (0.21) 0.69 (0.31)
Diff 0.00 (0.00) 1.00 (0.00) 0.93 (0.09) 1.00 (0.00)
Knowledge Larger_animal 0.72 (0.02) 0.63 (0.07) 0.81 (0.09) 0.84 (0.07)
Periodic_elements 0.99 (0.01) 0.96 (0.03) 1.00 (0.00) 0.98 (0.00)
Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.52 (0.09) 0.55 (0.11) 0.69 (0.15)
Common_concept 0.03 (0.02) 0.14 (0.04) 0.09 (0.04) 0.19 (0.05)
Object_counting 0.30 (0.02) 0.38 (0.06) 0.40 (0.12) 0.41 (0.03)
Odd_one_out 0.32 (0.02) 0.57 (0.02) 0.25 (0.18) 0.64 (0.00)
Orthography_starts_with 0.23 (0.01) 0.41 (0.09) 0.54 (0.06) 0.60 (0.12)
Taxonomy_animal 0.02 (0.02) 0.67 (0.14) 0.85 (0.06) 0.71 (0.02)
Auto_categorization 0.31 (0.01) 0.29 (0.02) 0.07 (0.07) 0.29 (0.04)
Word_sorting 0.58 (0.01) 0.64 (0.05) 0.23 (0.20) 0.70 (0.03)
CLUE Sentence_similarity 0.00 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.07)
Sentiment 0.90 (0.00) 0.88 (0.03) 0.88 (0.02) 0.89 (0.01)
Translation Num_to_verbal 0.13 (0.02) 0.99 (0.01) 1.00 (0.00) 0.99 (0.01)
Translation_en-de 0.83 (0.01) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01)
Translation_en-es 0.86 (0.01) 0.67 (0.24) 0.89 (0.00) 0.87 (0.02)
Translation_en-fr 0.88 (0.01) 0.77 (0.06) 0.85 (0.02) 0.83 (0.01)
Style Informal_to_formal 0.57 (0.01) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.25 (0.00)
median score 0.49 0.66 0.64 0.71
# best-performing tasks 9 8 7 14
Table 2: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, and INSTINCT. The bottom rows report the median score and total number of best-performing tasks for each method.
@@ 5 Experiments
We tackle instruction learning for ChatGPT. We conduct instruction induction tasks using 30 datasets from Honovich et al. (2023); Chen et al. (2024), along with a data summarization task from Gliwa et al. (2019). Furthermore, we compare the best-learned instructions from our ACING algorithm against human-provided instructions from Honovich et al. (2023), as shown in Table 1, with additional best-learned instructions in Table 9. Moreover, we compare ACING with three recent representative baselines for black-box instruction learning: APE (Zhou et al., 2023), InstructZero (Chen et al., 2024), and INSTINCT Lin et al. (2024), with results provided in Table 2.
Following previous works and to ensure fairness, we do not fine-tune the budget. Instead, like previous works, we consider a fixed budget of black-box LLM evaluations, set to $T=165$ Chen et al. (2024); Lin et al. (2024). Moreover, like previous works, we use Vicuna-13B Chiang et al. (2023) as the white-box LLM $h$. Furthermore, for previous works, we use their default (tuned) hyperparameters, including the intrinsic dimension $d^{\prime}=10$ and the number of soft tokens $N_{z}=5$. For fairness, we refrain from fine-tuning these parameters for our method and use the same values as in prior works. This ensures that our ACING algorithm searches in the same space, $[0,1]^{10}$, and uses the same total number of queries to the black-box LLM as APE, InstructZero, and INSTINCT for a fair comparison. For each algorithm, after identifying the best instruction using the validation set $\mathcal{V}$, we evaluate the discovered instruction on a separate test set $\mathcal{T}$ and report the test accuracy as the score.
We further conduct ablation studies on the key design choices for ACING, including the use of small learning budgets, providing reward plots for various selected tasks. Furthermore, we test the impact of using different white-box language models, intrinsic dimensions, and numbers of exemplars. More details on the experiments are provided in the Appendix.
@@
We use human instructions provided in Honovich et al. (2023) for various instruction-induction tasks, test them, and compare their scores against those of our best instructions. The complete list of selected tasks is provided in Table 1 in the Appendix. Several instructions were solved by both human experts and our approach, i.e., achieving a perfect test score of 1.00. For clarity and brevity, we omit these results from the main paper and focus only on tasks where the scores of humans and our method differed, as shown in Table 1.
As shown in these tables, our approach outperforms human instructions in several tasks, sometimes by a significant margin. For example, in the rhyming task, our method achieved a 0.39 improvement in the test score. The human instruction for this task—Write a word that rhymes with the input word.—only resulted in a score of 0.61. In contrast, the best ACING instruction, which was an intriguing prompt: Input the word that the program thought I was inputting and then output the word that program thought I was inputting, achieved a perfect score of 1.00.
Furthermore, in the challenging sentence similarity task, while the human-provided instruction achieved a score of 0.00, our approach improved the test score by up to 0.21. However, the instruction yielding the highest score appeared to bias the prediction towards outputting “3” due to the following instruction: Find a sentence pair that is probably not similar, and the output is 3 - probably, which resulted in the best score. The second-best instruction, proposed by ACING, achieved an improvement of 0.14 over the human score and was as follows: Generate a probability score for each pair of sentences, where 1 - probably not means that the two sentences are not related, 0 - definitely not means that the two sentences are not related, and 3 - probably means that the two sentences are related. Overall, our approach recovers several human test scores while being fully automatic, and in many cases surpasses human performance in terms of the number of successful tasks, average score, and median score across tasks. For the curious reader, we provide all our best instructions along with their test scores in Table 10 in the Appendix.
@@
We compare our method against recent baselines on the 30 instruction-induction datasets. Table 2 presents the average test accuracy (along with the standard deviation) over three independent runs, using three different seeds. For each seed, we selected the best instruction achieved by each method and evaluated it on the testing dataset. The table demonstrates that our method, ACING, outperforms others by achieving the highest accuracy in 14 out of 30 tasks, compared to INSTINCT, InstructZERO, and APE which succeeded in 9 tasks or less each. Additionally, ACING achieves the highest median accuracy across tasks, with a value of 0.71, which is  22 percentage points higher than APE. Further details, including the best prompt achieved for each task and corresponding test scores, are in the Appendix, Table 10. Moreover, we compare the performance of ACING against other methods on a summarization task using SAMSum dataset Gliwa et al. (2019) in Table 3, which shows that ACING outperforms the other methods.
In the following section, we demonstrate that the median score of ACING can be further improved by 5 percentage points using a simple two-phase exploration technique (splitting the budget), resulting in a 10-percentage-point improvement over the best baseline considered. Furthermore, in ablation studies, we show that fine-tuning the intrinsic dimensions $d^{{}^{\prime}}$ can lead to additional improvements.
Metric APE InstructZero INSTINCT ACING
ROUGE-1 0.35 (0.01) 0.33 (0.00) 0.36 (0.01) 0.37 (0.01)
ROUGE-2 0.12 (0.00) 0.11 (0.00) 0.14 (0.00) 0.14 (0.00)
ROUGE-L 0.25 (0.00) 0.24 (0.01) 0.27 (0.01) 0.28 (0.01)
@@
Table 3: Average test performance (and standard deviations) for instruction optimization for summarization task using SAMSum dataset.
5.3 Splitting the Budget
Due to the stochastic nature of the black-box LLM, the same instruction may yield different rewards when evaluated by the LLM. For more robust prompt selection, the budget $T$ can be split into two parts: an actor-critic exploration phase (repeating steps 1 to 4), followed by a uniform exploration phase. In the latter phase, the top $p$ prompts discovered by the actor-critic are evaluated multiple times, $k$ times each, using the black-box LLM. Specifically, the first phase uses $T-p\cdot k$ API calls, while the second uses the remaining calls. The prompt with the highest average reward across repetitions can then be used at test time. In Table 5 in the Appendix, we show that ACING, using this budget split technique with an initial exploration budget of $150$ and a remaining budget of $15$ calls (allocated to uniformly explore the top $p=5$ prompts, each evaluated $k=3$ times), achieves improved median scores across tasks, resulting in 10 percentage points improvement over the best baseline considered and higher test accuracy on 15 tasks compared to other methods.

### v2

@@ 3.1 Actor-Critic for Instruction Optimization
$\hat{\nabla}_{\mathbf{w}}J_{Q}(\mathbf{w})=\nabla_{\mathbf{w}}Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)\left(Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)-r\left(\mathbf{a}_{t}\right)\right).$ (5)
Training alternates between actor and critic updates. Despite the lack of state transitions, this architecture proves effective: the critic generalizes reward signals in prompt space, stabilizes updates by reducing reward variance, and enhances robustness via a twin-critic setup to mitigate overestimation. Empirically, this actor-critic formulation consistently outperforms no-critic baselines.
3.2 Enhancing Exploration via Entropy Regularization
In our setting, the agent operates under a fixed evaluation budget of $T$ queries and must discover the best-performing instruction within this limit. This corresponds to a pure exploration regime, where broad coverage of the action space is essential. While random exploration may seem sufficient, effective learning often requires balancing exploration with exploitation of prior observations to guide the search toward high-reward regions.
To encourage systematic exploration while leveraging past experience, we adopt the maximum entropy reinforcement learning framework Ziebart (2010); Haarnoja et al. (2018). This approach augments the expected reward with a policy entropy term, promoting stochasticity in the actor’s decisions. The objective becomes:
@@ 3.2 Enhancing Exploration via Entropy Regularization
where $\alpha$ is a temperature coefficient that governs the exploration-exploitation trade-off. A higher $\alpha$ encourages greater policy entropy, while a lower value biases the policy toward exploitation.
This formulation simplifies the soft actor-critic objective Haarnoja et al. (2018), omitting state-related components and long-horizon returns, since our environment is stateless and rewards are immediate. The entropy term, $\mathbb{E}_{\mathbf{a}\sim\pi(\cdot;\theta)}[\log\pi(\mathbf{a})]$, incentivizes the actor to maintain a diverse action distribution, thereby avoiding premature convergence to suboptimal prompts.
To avoid manual tuning of $\alpha$, we follow prior work Haarnoja et al. (2019) and adapt it to match a target entropy $H_{\text{target}}$, by minimizing:
@@
We use stochastic gradient descent, specifically the Adam optimizer Kingma and Ba (2017); Reddi et al. (2018), to jointly update the policy, critic, and entropy temperature.
3.3 Putting it All Together
Fig. 1 illustrates our actor-critic framework and its interaction with the environment. Fig. 2 zooms in on an example from the larger_animal dataset.
Overview: In each iteration $t\leq T$, the actor-critic agent generates a continuous vector “action” $\mathbf{a}$ (step 1). The action is then projected into the appropriate space using a fixed matrix $P$ to obtain $\mathbf{z}$. The environment then concatenates the projected vector $\mathbf{z}$ with a set of exemplars’ embeddings from $\mathcal{E}$ and feeds it into a white-box model $h$ (step 2). The white-box model produces a discrete prompt, $\tau$, which is evaluated using the validation dataset $\mathcal{V}$ based on the responses from the black-box LLM $f$ (step 3). The black-box LLM’s prediction is compared to true labels of the validation examples, and a score function provides a reward, used to update both the critic and actor networks accordingly.
@@ 3.3 Putting it All Together
After $T$ iterations, the agent returns the best-performing prompt $\tau^{\star}$, which is evaluated on the test set $\mathcal{T}$ in the black-box LLM using the evaluation template shown in Fig. 1 (bottom right).
Category Task Human Instruction ACING Instruction (Ours) Human ACING
Spelling Second_word_letter Extract the second letter of the input word. Input a word and output the letter that corresponds to the second letter in that word 0.96 (0.00) 0.92 (0.00)
Syntax Negation Negate the input sentence. Flip the truth value of the statements in the input 0.81 (0.00) 0.82 (0.00)
Lexical Semantics Antonyms Write a word that means the opposite of the input word. Take a word and change it to its opposite 0.70 (0.00) 0.83 (0.00)
Synonyms Write a word with a similar meaning to the input word. Input a word that is a synonym for the word that was output 0.14 (0.01) 0.13 (0.00)
Phonetics Rhymes Write a word that rhymes with the input word. Input the word that the program thought I was inputting and then output the word that program thought I was inputting 0.61 (0.01) 1.00 (0.00)
Semantics Cause_and_effect Find which of the two given cause and effect sentences is the cause. Find the sentence that is the cause of the effect in the pair of sentences 0.97 (0.02) 0.90 (0.02)
Style Informal_to_formal Rephrase the sentence in formal language. Convert the input into output using the same word order and with the same meaning 0.63 (0.00) 0.50 (0.00)
Multi- lingual Translation_en-de Translate the word into German. Provide a translation for each word in the English text into German 0.81 (0.00) 0.84 (0.00)
Translation_en-es Translate the word into Spanish. Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 0.89 (0.00) 0.88 (0.00)
Translation_en-fr Translate the word into French. Create a program that would take an English word as input and output its French equivalent 0.86 (0.00) 0.87 (0.00)
GLUE Sentiment Determine whether a movie review is positive or negative. Classify each input as positive or negative based on the assessment of the corresponding movie 0.89 (0.01) 0.91 (0.00)
Sentence_similarity Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly. Find a sentence pair that is probably not similar, and the output is 3 - probably 0.00 (0.00) 0.21 (0.00)
median score 0.81 0.86
@@
Table 2: Tasks from the instruction-induction datasets where the human and ACING test scores differed. For each task, we provide the corresponding human instruction as proposed in Honovich et al. (2023) and our best discovered instruction. We tested these instructions on the test dataset and report the average score (with standard deviation) over 3 repetitions.
Category Task APE EvoPrompt InstructZero INSTINCT ACING
Spelling Letters_list 0.59 (0.02) 0.97 (0.03) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00)
Second_word_letter 0.00 (0.00) 0.63 (0.17) 0.35 (0.09) 0.39 (0.28) 0.70 (0.15)
Morpho-Syntax Negation 0.79 (0.00) 0.84 (0.02) 0.65 (0.10) 0.58 (0.22) 0.71 (0.06)
Lexical Semantics Synonyms 0.14 (0.01) 0.19 (0.07) 0.22 (0.11) 0.19 (0.08) 0.13 (0.02)
Word_unscrambling 0.54 (0.00) 0.44 (0.06) 0.59 (0.06) 0.54 (0.02) 0.50 (0.07)
Phonetics Rhymes 0.59 (0.01) 0.52 (0.05) 0.99 (0.01) 0.36 (0.04) 0.57 (0.31)
Numerical Sum 0.87 (0.01) 1.00 (0.00) 1.00 (0.00) 0.70 (0.21) 1.00 (0.00)
Diff 0.00 (0.00) 0.99 (0.01) 1.00 (0.00) 0.93 (0.09) 1.00 (0.00)
Knowledge Larger_animal 0.72 (0.02) 0.58 (0.06) 0.63 (0.07) 0.81 (0.09) 0.84 (0.07)
Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.48 (0.10) 0.52 (0.09) 0.55 (0.11) 0.69 (0.15)
Common_concept 0.03 (0.02) 0.17 (0.00) 0.14 (0.04) 0.09 (0.04) 0.19 (0.05)
Object_counting 0.30 (0.02) 0.50 (0.06) 0.38 (0.06) 0.40 (0.12) 0.41 (0.03)
Odd_one_out 0.32 (0.02) 0.64 (0.04) 0.57 (0.02) 0.25 (0.18) 0.64 (0.00)
Orthography_starts_with 0.23 (0.01) 0.47 (0.02) 0.41 (0.09) 0.54 (0.06) 0.60 (0.12)
Taxonomy_animal 0.02 (0.02) 0.38 (0.15) 0.67 (0.14) 0.85 (0.06) 0.71 (0.02)
Auto_categorization 0.31 (0.01) 0.20 (0.03) 0.29 (0.02) 0.07 (0.07) 0.29 (0.04)
Word_sorting 0.58 (0.01) 0.01 (0.00) 0.64 (0.05) 0.23 (0.20) 0.70 (0.03)
CLUE Sentence_similarity 0.00 (0.00) 0.05 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.07)
Translation Num_to_verbal 0.13 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 0.99 (0.01)
Translation_en-es 0.86 (0.01) 0.76 (0.00) 0.67 (0.24) 0.89 (0.00) 0.87 (0.02)
Style Informal_to_formal 0.57 (0.01) 0.50 (0.02) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.25 (0.00)
median score 0.31 0.50 0.59 0.54 0.69
# best-performing tasks 3 7 8 4 13
Table 3: Average test performance (with standard deviations) over 3 seeds comparing ACING to APE (Zhou et al., 2023), EvoPrompt (Guo et al., 2024), InstructZero (Chen et al., 2024), and INSTINCT (Lin et al., 2024b) on the 23 most challenging tasks (where at least one method has score < 0.7). Bottom rows show median scores and the number of best-performing tasks.
@@ 4 Experiments
We focus on instruction learning for ChatGPT OpenAI (2023b), with additional analysis on GPT-4 OpenAI (2023c) and GPT-4o Hurst et al. (2024) as representative black-box LLMs. We conduct instruction induction tasks using 30 datasets spanning several diverse categories from Honovich et al. (2023); Chen et al. (2024), zero-shot CoT reasoning datasets (GSM8K Cobbe et al. (2021), AQUARAT Ling et al. (2017)), and summarization on SAMSum Gliwa et al. (2019).
In §4.1, we compare ACING’s best-learned instructions against human-written prompts from Honovich et al. (2023). We also benchmark, in §4.2, against four recent black-box instruction optimization methods: APE Zhou et al. (2023), EvoPrompt Guo et al. (2024), InstructZero Chen et al. (2024), and INSTINCT Lin et al. (2024b). Furthermore, we study the interpretability and clarity of generated instructions (§4.3) and conduct ablation studies (§4.4) to examine the impact of the different design choices.
To ensure fairness and comparability with other automatic approaches, we adopt a fixed evaluation budget of black-box queries T=165T=165, [^1] See Appendix F.3 for reward plots where ACING often peaks well before the budget is exhausted. consistent with prior work. Moroever, ACING uses an off-the-shelf, publicly available white-box decoder to map latent vectors to prompts. Moreover, all learning and evaluation occur solely through black-box interactions, which is consistent with other leading methods (e.g., INSTINCT, InstructZero). To ensure fairness, we fix the same decoder Vicuna-13B Chiang et al. (2023), as used in their analyses across all methods, while in ablations we compare with WizardLM-13B Xu et al. (2024).
The experimental setup isolates ACING’s core contribution, ensuring that performance gains are not attributable to decoder selection or tuning. Further details, including hyperparameters, are provided in Appendix B.
@@
We compare instructions found by ACING against human-authored prompts from Honovich et al. (2023) across a broad range of instruction-induction tasks. Table 2 highlights only those tasks where test performance differed between the two. The full set of tasks is included in Appendix G.
Acing not only matches but often surpasses human-written instructions—often by substantial margins. For instance, in the Antonyms task, the human instruction (“Write a word that means the opposite of the input word”) scores 0.70. ACING improves this to 0.82 with a more direct and actionable phrasing: “Take a word and change it to its opposite.” The formulation is simpler yet effective.
Consider the rhyming task, where the human instruction—“Write a word that rhymes with the input word”—yields a score of 0.61. ACING significantly improves performance, achieving a perfect score of 1.00 with an alternative phrasing. This highlights ACING’s ability to discover high-performing instructions that align closely with the underlying model behavior.
In the sentence similarity task, where the human instruction results in a score of 0.00, ACING raises performance to 0.21. The highest-scoring ACING instruction introduces a mild incline toward a mid-scale output (“3 - probably”), but its phrasing remains semantically valid and interpretable. A second-best instruction, unbiased, still improves upon the human-written version by 0.14 points.
On average, Acing improves the median task score from 0.81 to 0.86 and outperforms human-written instructions on 7 out of 12 tasks in this subset—underscoring its potential as a practical and effective alternative to manual prompt engineering. Further generated instructions are in Appendix I. Furthermore, extended results with GPT-4o are provided in Table˜11 in Appendix˜E.
@@
Instruction-induction datasets: In Table 3, we show tasks from Honovich et al. (2023) where at least one method failed to achieve 70% score (full results on all 30 tasks can be found in Appendix˜C). The table shows the average test accuracy (along with the standard deviation) over three independent runs, using three different seeds. For each seed, we selected the best instruction achieved by each method and evaluated it on the testing dataset. The results demonstrate that our method, ACING, outperforms the others, achieving the highest accuracy in 13 out of the remaining 23 tasks, compared to INSTINCT, InstructZERO, EvoPrompt, and APE, which succeeded in 8 tasks or fewer. Additionally, ACING achieves the highest median accuracy across tasks, with a value of 0.69, which is approximately 10 percentage points higher than the best baseline. The score types can be found in Table 8. Moreover, the best prompt achieved for each task and the corresponding test scores, can be found in Table 21 in Appendix˜C. Further analyses on GPT-4 can be found in the Appendix. D.
CoT datasets. We evaluate our method on two zero-shot CoT reasoning datasets: GSM8K Cobbe et al. (2021) and AQUA-RAT Ling et al. (2017). Prior work Kojima et al. (2022) shows that simple CoT prompts can enhance LLM performance. Using 100 steps (and other settings as above), ACING achieves the best results on both, demonstrating strong CoT reasoning capability.
Summarization dataset. We compare the performance of ACING with other methods on summarization tasks using the SAMSum dataset Gliwa et al. (2019). The results, presented in Table 4, show that ACING outperforms the other methods across the three metrics considered: ROUGE-1, ROUGE-2, and ROUGE-L Lin (2004).
Statistical significance test. We conduct a Wilcoxon signed-rank test Wilcoxon (1992), a non-parametric test. The results confirm that ACING significantly outperforms all baselines across tasks, with p = 0.0005 (APE), 0.0041 (INSTINCT), 0.0092 (EvoPrompt), and 0.0335 (InstructZero), all below the standard 0.05 threshold, indicating that the observed gains are statistically significant.
Metric APE EvoPompt InstructZero INSTINCT ACING
ROUGE-1 0.35 (0.01) 0.35 (0.01) 0.33 (0.00) 0.36 (0.01) 0.37 (0.01)
ROUGE-2 0.12 (0.00) 0.12 (0.00) 0.11 (0.00) 0.14 (0.00) 0.14 (0.00)
ROUGE-L 0.25 (0.00) 0.26 (0.00) 0.24 (0.01) 0.27 (0.01) 0.28 (0.01)
@@
Table 4: Average test performance (and standard deviations) for summarization task using SAMSum dataset.
Results breakdown. To better understand ACING’s strengths, we grouped the tasks into various categories. On cognitive tasks, requiring complex reasoning, ACING demonstrates the clearest advantage. It ranks first on 5 out of 8 such tasks from Table 3, including Cause_and_effect (0.69, +14%), Word_sorting (0.70, +6%), Odd_one_out (0.64, tied), and Orthography_starts_with (0.60, +6%). As shown in Fig. 3 (ordered left to right by median ranking), ACING achieves the best median rank and shows a strong skew toward top rankings in this category. Furthermore, ACING shows the best performance in the zero-shot CoT reasoning tasks (Table˜5). In symbolic manipulation, ACING again performs best, achieving perfect scores on Letters_list and First_word_letter, and leading on Second_word_letter (0.70 vs. 0.63). This highlights its precision in low-level, structured tasks. ACING is competitive but not dominant in tasks involving lexical semantics, world knowledge, and translation—e.g., Informal_to_formal (0.44 vs. 0.57), and Translation_en-es (0.87 vs. 0.89). Finally, it consistently outperforms the other methods in summarization.
Method Dataset Best Zero-Shot Instruction Score
Kojima et al. (2022) GSM8K Let’s think step by step. 0.72
INSTINCT Lin et al. (2024b) GSM8K Let’s use our creativity to find the solution. 0.75
ACING (Ours) GSM8K Let’s use our math skills to conquer this challenge. 0.76
Kojima et al. (2022) AQUA-RAT Let’s think step by step. 0.59
INSTINCT Lin et al. (2024b) AQUA-RAT Let’s break it down. 0.59
ACING (Ours) AQUA-RAT Let’s use the power of substitution to solve this problem. 0.63
Table 5: Best zero-shot instructions and corresponding scores for different methods and datasets.
4.3 Instruction Clarity and Readability
To evaluate the interpretability and clarity of generated instructions, we assess whether ACING-generated prompts are understandable and task-aligned. We conduct a human evaluation with 26 participants who rated the clarity and alignment of generated instructions from Table˜2 on a 5-point Likert scale. The results show that participants found the prompts generally clear, with a median score of 3.9/5 (see Appendix H for the protocol and statistics). In parallel, we use standard readability metrics to automatically assess the generated prompts. The results show a median Flesch Reading Ease (FRE) of 70.8, Flesch-Kincaid Grade Level (FKG) of 7.0, and Coleman-Liau Index (CLI) of 7.3. These scores correspond to mid-grade readability (7th–8th grade), indicating that the generated instructions are accessible to a broad user base.